TOOP TREIBle answers in the space provided. 1 Whst is the shugle of a projectile launcheed fron the ground which has a maxinum height that is hals its maximum range?

media%2F134%2F1340f9b2-198b-44b7-b910-6c
Answer #1

answer 1) the maximum range is given by

Rmax=1/2 v2sin2\Theta/g............1)

and maximum height is given by

H=v2sin2\Theta/2g......2)

equating 1 and 2 we have

v2sin2\Theta/2g=1/2*v2sin2\Theta/g

2sin2\Theta/2=2sin\Thetacos\Theta

sin\Theta=2cos\Theta

tan\Theta=2

\Theta=tan-1(2)=63.4o

so the answer is 63.4o